# Relative monodromy of ramified sections on abelian schemes

### v3

@@ 7. $(vii)$
Let us point out the relationship between the cohomology groups $H^{1}(\pi_{1}(S^{\prime}(\mathbb{C}),s^{\prime}),\mathbb{Z}^{2g})$ and $H^{1}(\pi_{1}(C^{\prime},s^{\prime}),\mathbb{Z}^{2g})$. By part $(iii)$ of 3.6 the homomorphism $i^{\prime}_{*}:\pi_{1}(C^{\prime},s^{\prime})\to\pi_{1}(S^{\prime},s^{\prime})$ is surjective and this induces a homomorphism
@@
Let $C$ be a curve satisfying 3.6. We work by contradiction by assuming $M_{\sigma}^{\textrm{rel}}=\{0\}$. By part $(vii)$ of 3.6 this implies $M_{\sigma_{C}}^{\textrm{rel}}=\{0\}$, or equivalently $\ker{\rho_{C}}\subseteq\ker{\theta_{\sigma_{C}}}$. Let $q_{C*}:\pi_{1}(C,s)\to\pi_{1}(C^{\prime},s^{\prime})$ be the homomorphism appearing in Equation 24. By part $(v)$ of 3.6 we have $\ker{q_{C*}}\subseteq\ker{\rho_{C}}$; thus, we have $\ker{q_{C*}}\subseteq\ker{\theta_{\sigma_{C}}}$. Introducing the notation $\overline{G}:=q_{C*}\pi_{1}(C,s)$ we obtain that there exists a homomorphism $f:\overline{G}\rightarrow\textrm{SL}_{2g+1}(\mathbb{Z})$ such that the diagram
@@
commutes; in other words, $\theta_{\sigma_{C}}$ factors through $q_{C*}$. By using the same notation as in Equation 13, this means that for $h\in\pi_{1}(C,s)$ the vector $c(h)$ only depends on the element $\overline{h}:=q_{C*}(h)\in\overline{G}$. Then we obtain a cocycle
@@
To this end, let us construct some homomorphisms between cohomology groups. Since we are assuming $M_{\sigma}^{\textrm{rel}}=\{0\}$ and the map $q_{*}:\pi_{1}(S(\mathbb{C}),s)\to\pi_{1}(S^{\prime}(\mathbb{C}),s^{\prime})$ is surjective by 3.4, with the same reasoning as in Equation 29 we obtain an injective homomorphism
@@
which to a cohomology class $[\tau]_{S}$ associates a cohomology class denoted with $[\tau]_{S^{\prime}}$. By 3.10 we have a homomorphism
@@
where $\tau\in\Sigma(S)$ is such that $r(\tau)=\tau_{C}$: notice that this map is well-defined by 3.6. In fact, if $\tau_{1},\tau_{2}\in\Sigma(S)$ are such that $\tau_{1,C}=\tau_{2,C}$, then $M_{\tau_{1}}^{\textrm{rel}}=M_{\tau_{2}}^{\textrm{rel}}$ by part $(vi)$ of 3.6. This implies $[\tau_{1}]_{S}=[\tau_{2}]_{S}$ and so $[\tau_{1}]_{C^{\prime}}=[\tau_{2}]_{C^{\prime}}$.
@@
$\xi|_{r(H^{0}(S,\Sigma^{\operatorname{an}}))}:r(H^{0}(S,\Sigma^{\operatorname{an}}))\to H^{1}(\overline{G},\mathbb{Z}^{2g}),\qquad\tau_{C}\mapsto[\tau_{C}]_{\overline{G}}.$
@@
$\eta^{\prime}:=\eta|_{\textrm{Im}(\zeta)}:\textrm{Im}(\zeta)\to\textrm{Im}(\xi|_{r(H^{0}(S,\Sigma^{\operatorname{an}}))}),\qquad[\tau]_{C^{\prime}}\mapsto[\tau_{C}]_{\overline{G}},$
where $r(\tau)=\tau_{C}$; let’s prove that $\eta^{\prime}$ is an isomorphism. Notice that it is well-defined and surjective by construction, so it’s enough to prove that it is injective. To this end, consider $\tau_{1},\tau_{2}\in\Sigma(S)$ such that $[\tau_{1,C}]_{\overline{G}}=[\tau_{2,C}]_{\overline{G}}$. This implies
@@
Thanks to the injective homomorphism obtained by Equation 30, we get $[\tau_{1,C}-\tau_{2,C}]_{C}=0$. By part $(vi)$ of 3.6, we also obtain $[\tau_{1}-\tau_{2}]_{S}=0$. By applying $\zeta_{3}$ we get $[\tau_{1}-\tau_{2}]_{C^{\prime}}=0$, which means that $\eta^{\prime}$ is injective.
Thus, by what we have just proven and by 3.9 the section $\sigma_{C}$ determines a cohomology class $[\sigma_{C}]_{C^{\prime}}$ in $H^{1}(\pi_{1}(C^{\prime},s^{\prime}),\mathbb{Z}^{2g})\cong H^{1}(C^{\prime},\Lambda_{C^{\prime}})$. Again thanks to 3.9, we have that there exists a (possibly transcendental) section $\sigma^{\prime}:C^{\prime}\rightarrow\mathcal{A}^{\prime}|_{C^{\prime}}$ whose cohomology class $[\sigma^{\prime}]$ in $H^{1}(\pi_{1}(C^{\prime},s^{\prime}),\mathbb{Z}^{2g})$ coincides with the class $[\sigma_{C}]_{C^{\prime}}$ defined above, i.e. $[\sigma^{\prime}]=[\sigma_{C}]_{C^{\prime}}$. Therefore, by 2.9 we obtain a section of the form $\sigma_{C}-q_{C}^{*}(\sigma^{\prime})\in\Sigma^{\operatorname{an}}(C)$ that admits a logarithm; let us write
@@
Now, by using properties of the operators $q_{C}^{*}$ and $\operatorname{Tr}$, we will get a contradiction. Applying first the trace operator and then the pullback operator we obtain:
@@
By part $(vi)$ of 3.6, the section $q^{*}\operatorname{Tr}(\sigma)-\deg{q}\cdot\sigma$ is also an algebraic section admitting a global logarithm on $S(\mathbb{C})$. By 3.3 and [22, Proposition 2.10] the base $S$ can be thought to be maximal. By 2.8 we must have $\sigma=q^{*}\operatorname{Tr}(\sigma)/\deg{q}$, which leads to a contradiction with the fact that $\sigma$ is a ramified section.
Remark 3.11.
Note that, under the hypothesis of generic finiteness of the modular map, from 2.6 and 1.2 it follows immediately that a torsion section $\sigma\colon S\to\mathcal{A}$ is unramified. When $S$ is a curve this follows also from [31, Theorem 1].
@@ 3.3 Rank of relative monodromy
Remark 3.12.
When $\mathcal{A}\to S$ is simple and $S$ is a curve, maximal variation in moduli is equivalent to not having fixed part. Hence our 1.2 holds for all ramified sections of an abelian scheme with no fixed part over a curve $S$.
An interesting case where our result holds, is given by the family of hyperelliptic curves described in [3, Section 7].
The hypothesis on the irreducibily of the action of $\text{Mon}(\mathcal{A})$ appearing in 1.3 is satisfied when $p(S)$ is not contained in any proper special subvariety of $\mathbb{A}_{g}$.
3.4 Relative monodromy of non-torsion sections
So far all the results are proved for ramified sections, which are a subset of all non-torsion sections (see 3.11). We can extend our main results to all non-torsion sections under the assumption that the modular map $p$ is generically finite onto a subvariety $\mathcal{G}\subseteq\mathbb{A}_{g}$, where the unramified base changes of $\mathcal{G}$ have finite Mordel-Weil group.
Proposition 3.13.
Let $\phi:\mathcal{A}\to S$ be an abelian scheme without fixed part such that (up to a finite base change) the modular map $p:S\to\mathcal{G}\subseteq\mathbb{A}_{g}$ is generically finite, where the unramified base changes of $\mathcal{G}$ have finite Mordel-Weil group. If $\sigma:S\to\mathcal{A}$ is a non-torsion section, then $M_{\sigma}^{\textrm{rel}}$ is non-trivial. If, in addition, the action of the monodromy group $\textrm{Mon}(\mathcal{A})$ is irreducible, then $M_{\sigma}^{\textrm{rel}}$ is isomorphic to $\mathbb{Z}^{2g}$.
@@ 3.4 Relative monodromy of non-torsion sections
Since the unramified base changes of $\mathcal{G}$ have finite Mordell-Weil group, each unramified section is torsion. The result immediately follows from 1.2 and 1.3.
The condition about the Mordell-Weil group holds true when $\mathcal{G}$ is a Kuga family (see [26, Main Theorem]). In particular, $\mathcal{G}$ may be chosen to be a totally geodesic (i.e. weakly-special) subvariety of $\mathbb{A}_{g}$. Notice that, if we pick $\mathcal{G}=\mathbb{A}_{g}$ we obtain the complete version of 1.4: in fact we don’t need to distinguish the cases when the image of the section $\sigma$ is contained in proper group subschemes since having a generically finite map on a universal family prevents our abelian scheme from having any non-trivial proper subscheme; for the same reason, in this case the scheme automatically does not have fixed part.
3.5 The case of thin Monodromy of periods
In [2] André uses some Hodge-theoretical methods to address the problems presented in this paper. These are his hypotheses:
  • $(i)$
    $\mathcal{A}\to S$ has no fixed part, even after pull-back to a finite étale cover of $S$,
  • $(ii)$
    the image of the section is not contained in any proper group-subgroup,
  • $(iii)$
    the monodromy group $\textrm{Mon}(\mathcal{A})$ is not thin, i.e. it is of finite index in the integral points of its Zariski closure,
He shows that $M_{\sigma}^{\textrm{rel}}\cong\mathbb{Z}^{2g}$. It would be interesting to compare his results with ours.
Regarding the hypothesis $(iii)$, André observes that it is not always satisfied and in general we don’t know how big is $M_{\sigma}^{\textrm{rel}}$ in such case. To this regard, we provide below an explicit construction showing a case where $M_{\sigma}^{\textrm{rel}}$ is of maximal rank even though $\textrm{Mon}(\mathcal{A})$ is thin.
Remark 3.14.
Notice that a priori, our hypotheses in 1.3 don’t exclude the case when $\textrm{Mon}(\mathcal{A})$ is thin.
Example 3.15.
Let us denote by $\mathcal{L}\to S$ the Legendre elliptic scheme, where $S:=\mathbb{P}_{1}-\{0,1,\infty\}$. Consider the product
$(\phi_{1},\phi_{2}):\mathcal{L}\times\mathcal{L}\to S\times S,$
where we use the coordinates $(x,y)$ in $S\times S$. We let $X$ be the line in $S^{2}$ defined by $x+y=2$, so this line is isomorphic under the first projection to $B:=\mathbb{P}_{1}-\{0,1,2,\infty\}$. By projection from $X$ to $B$, we may define a scheme $\phi:\mathcal{A}\to B$ over $B$ whose fibers are of the type $\phi^{-1}(\lambda)=\mathcal{L}_{\lambda}\times\mathcal{L}_{2-\lambda}$.
The abelian scheme $\phi:\mathcal{A}\to B$ may be interpreted as the fiber product of two elliptic schemes $\mathcal{E}_{1}\times_{B}\mathcal{E}_{2}\to B$, where $\mathcal{E}_{1}\to B$ is the Legendre scheme associating $\mathcal{L}_{\lambda}\to\lambda$ and $\mathcal{E}_{2}\to B$ is the Legendre scheme associating $\mathcal{L}_{\lambda}\to 2-\lambda$. Note that the two elliptic schemes are not isogenous, since they have a different set of bad reduction.
The fundamental group $G$ of $B(\mathbb{C})$ is free on three generators, and the monodromy representation on the periods of $\phi$ gives a homomorphism $\rho:G\to\Gamma_{2}\times\Gamma_{2}$, where $\Gamma_{2}$ is the subgroup of $\textrm{SL}_{2}(\mathbb{Z})$ generated by
$\begin{pmatrix}1&2\\ 0&1\end{pmatrix},\qquad\begin{pmatrix}1&0\\ 2&1\end{pmatrix}.$
We denote the monodromy group of periods as $\textrm{Mon}(\mathcal{A}):=\rho(G)$. Since $\phi:\mathcal{A}\to B$ is the product of two non-isogenous elliptic schemes, then the Zariski closure of $\textrm{Mon}(\mathcal{A})$ is the whole $\textrm{SL}_{2}\times\textrm{SL}_{2}$ [9, Theorem 5.3.10].
The abelian scheme $\phi\colon\mathcal{A}\to B$ has trivial Mordell-Weil rank. Let us consider the map $\sigma:B\to\mathcal{A}$ defined by the matrices
This map is a multisection of the abelian scheme $\phi$, but it becomes a section whether we consider the base change induced by the degree $4$ field extension $\mathbb{C}(B)\subseteq\mathbb{C}(B)(\psi,\mu)$ where $\psi^{2}+\mu^{2}=3$ and $\mu^{2}=1+\lambda$. Denote by $\phi^{\prime}:\mathcal{A}^{\prime}\to B^{\prime}$ the base change we just described. The monodromy of periods $\textrm{Mon}(\mathcal{A}^{\prime})$ is a finite index subgroup of $\textrm{Mon}(\mathcal{A})$: in particular, it is Zariski dense in $\textrm{SL}_{2}\times\textrm{SL}_{2}$ and it has infinite index in $\Gamma_{2}\times\Gamma_{2}$; so $\textrm{Mon}(\mathcal{A}^{\prime})$ is still thin. The morphism $\sigma$ is a section of $\phi^{\prime}$ whose image is not contained in any proper group subscheme, thus the relative monodromy group $M_{\sigma}^{\textrm{rel}}$ is isomorphic to $\mathbb{Z}^{4}$ by [33, Theorem 4.11].
4 Some applications
Now, we give some useful applications of our results.
4.1 Manin’s kernel theorem
A first application of the non-triviality of relative monodromy is a strong version of Manin’s kernel theorem under the hypotheses of our 3.13. In the case of elliptic schemes the same result was obtained in [10]. We refer to the Betti coordinates defined in Section 2.2 (see Equation 5).
@@ 4.1 Manin’s kernel theorem
Let $\phi:\mathcal{A}\to S$ be an abelian scheme without fixed part such that (up to a finite base change) the modular map $p:S\to\mathcal{G}\subseteq\mathbb{A}_{g}$ is generically finite, where the unramified base changes of $\mathcal{G}$ have finite Mordel-Weil group. Let $\sigma:S\to\mathcal{A}$ be a section. If the Betti coordinates $\beta_{\sigma,i}$ with $i=1,\ldots,2g$ are globally defined functions on $S(\mathbb{C})$, then $\sigma$ is a torsion section. In that case, the Betti coordinates are rational constants.
Proof.
If the Betti coordinates are globally defined on $S(\mathbb{C})$ then the relative monodromy group $M_{\sigma}^{\textrm{rel}}$ is trivial. The claim then follows by the non-triviality part of 3.13. ∎
4.2 Algebraic independence of periods and logarithms
Our theorem about the rank of the relative monodromy group 1.3 gives a different proof of a result due to André [1, Theorem 3] about the algebraic independence of logarithms under our hypothesis on the abelian scheme $\mathcal{A}\to S$. Let’s briefly describe the setting.
Recall that the monodromy action of the fundamental group $\pi_{1}(S(\mathbb{C}),s)$ on periods and the logarithm of a section $\sigma$ induces a representation $\theta_{\sigma}:\pi_{1}(S(\mathbb{C}),s)\rightarrow\textrm{SL}_{2g+1}(\mathbb{Z})$ described in Equation 13. Consider the projection $\theta_{\sigma}(\pi_{1}(S(\mathbb{C}),s))\to\textrm{Sp}_{2g}(\mathbb{Z})$, whose image is the monodromy group $\textrm{Mon}(\mathcal{A})$. By general theory, the Zariski closure $\overline{\theta_{\sigma}(\pi_{1}(S(\mathbb{C}),s))}$ in $\textrm{SL}_{2g+1}$ is the differential Galois group of the Picard-Vessiot extension of $\mathbb{C}(S)$ generated by the coordinates of $\omega_{1},\ldots,\omega_{2g},\log_{\sigma}$ and their directional derivatives along a tangent vector field $\partial$ with respect to the Gauss-Manin connection. Clearly the homomorphism $\theta_{\sigma}(\pi_{1}(S(\mathbb{C}),s))\to\textrm{Sp}_{2g}(\mathbb{Z})$ extends to an algebraic group homomorphism
$\xi:\overline{\theta_{\sigma}(\pi_{1}(S(\mathbb{C}),s))}\to\textrm{Sp}_{2g}$
Theorem 4.2.
Let $\phi:\mathcal{A}\to S$ be an abelian scheme without fixed part such that (up to a finite base change) the modular map $p:S\to p(S)\subseteq\mathbb{A}_{g}$ is generically finite. If $\sigma:S\to\mathcal{A}$ is a ramified section and the action of the monodromy group $\textrm{Mon}(\mathcal{A})$ is irreducible, then the kernel $\ker{\xi}$ is isomorphic to $\mathbb{G}_{a}^{2g}$. In particular, the coordinates of $\log_{\sigma}$ have transcendence degree $g$ over $\mathbb{C}(S)(\omega_{1},\ldots,\omega_{2g})$, where $\mathbb{C}(S)(\omega_{1},\ldots,\omega_{2g})$ denotes the extension of $\mathbb{C}(S)$ generated by the coordinates of periods.

### v4

@@ 7. $(vii)$
Let us point out the relationship between the cohomology groups $H^{1}(\pi_{1}(S^{\prime}(\mathbb{C}),s^{\prime}),\mathbb{Z}^{2g})$ and $H^{1}(\pi_{1}(C^{\prime},s^{\prime}),\mathbb{Z}^{2g})$. By part $(iii)$ of 4.5 the homomorphism $i^{\prime}_{*}:\pi_{1}(C^{\prime},s^{\prime})\to\pi_{1}(S^{\prime},s^{\prime})$ is surjective and this induces a homomorphism
@@
Let $C$ be a curve satisfying 4.5. We assume $M_{\sigma}^{\textrm{rel}}=\{0\}$ and we prove that $\sigma$ is unramified over $S$. By part $(vii)$ of 4.5 this implies $M_{\sigma_{C}}^{\textrm{rel}}=\{0\}$, or equivalently $\ker{\rho_{C}}\subseteq\ker{\theta_{\sigma_{C}}}$. Let ${q_{C}}_{\ast}:\pi_{1}(C,s)\to\pi_{1}(C^{\prime},s^{\prime})$ be the homomorphism appearing in Equation 28. By part $(v)$ of 4.5 we have $\ker{{q_{C}}_{\ast}}\subseteq\ker{\rho_{C}}$; thus, we have $\ker{{q_{C}}_{\ast}}\subseteq\ker{\theta_{\sigma_{C}}}$. Introducing the notation $\overline{G}:={q_{C}}_{\ast}\pi_{1}(C,s)$ we obtain that there exists a homomorphism $f:\overline{G}\rightarrow\textrm{SL}_{2g+1}(\mathbb{Z})$ such that the diagram
@@
commutes; in other words, $\theta_{\sigma_{C}}$ factors through ${q_{C}}_{\ast}$. By using the same notation as in Equation 16, this means that for $h\in\pi_{1}(C,s)$ the vector $c(h)$ only depends on the element $\overline{h}:={q_{C}}_{\ast}(h)\in\overline{G}$. Then we obtain a cocycle
@@
To this end, let us construct some homomorphisms between cohomology groups. Since we are assuming $M_{\sigma}^{\textrm{rel}}=\{0\}$ and the map $q_{*}:\pi_{1}(S(\mathbb{C}),s)\to\pi_{1}(S^{\prime}(\mathbb{C}),s^{\prime})$ is surjective by 4.4, with the same reasoning as in Equation 31 we obtain an injective homomorphism
@@
which to a cohomology class $[\tau]_{S}$ associates a cohomology class denoted with $[\tau]_{S^{\prime}}$. By 4.8 we have a homomorphism
@@
where $\tau\in\Sigma(S)$ is such that $r(\tau)=\tau_{C}$: notice that this map is well-defined by 4.5. In fact, if $\tau_{1},\tau_{2}\in\Sigma(S)$ are such that $\tau_{1,C}=\tau_{2,C}$, then $\theta_{\tau_{1}}=\theta_{\tau_{2}}$ by part $(iii)$ and $(vi)$ of 4.5. This implies $[\tau_{1}]_{S}=[\tau_{2}]_{S}$ and so $[\tau_{1}]_{C^{\prime}}=[\tau_{2}]_{C^{\prime}}$.
@@
$\xi^{\prime}:=\xi|_{r(H^{0}(S,\Sigma^{\operatorname{an}}))}:r(H^{0}(S,\Sigma^{\operatorname{an}}))\to H^{1}(\overline{G},\mathbb{Z}^{2g}),\qquad\tau_{C}\mapsto[\tau_{C}]_{\overline{G}}.$
@@
$\eta^{\prime}:=\eta|_{\textrm{Im}(\zeta)}:\textrm{Im}(\zeta)\to H^{1}(\overline{G},\mathbb{Z}^{2g}),\qquad[\tau]_{C^{\prime}}\mapsto[\tau_{C}]_{\overline{G}},$
where $r(\tau)=\tau_{C}$. Thanks to Equation 28, we have $\textrm{Im}(\eta^{\prime})=\textrm{Im}(\xi^{\prime})$ and the following commutative diagram:
[…]
Let’s prove that $\eta^{\prime}$ is an isomorphism. Notice that it is well-defined and surjective by construction, so it’s enough to prove that it is injective. To this end, consider $\tau_{1},\tau_{2}\in\Sigma(S)$ such that $[\tau_{1,C}]_{\overline{G}}=[\tau_{2,C}]_{\overline{G}}$. This implies
@@
Thanks to the injective homomorphism obtained by Equation 32, we get $[\tau_{1,C}-\tau_{2,C}]_{C}=0$. By part $(vi)$ of 4.5, we also obtain $[\tau_{1}-\tau_{2}]_{S}=0$. By applying $\zeta_{3}$ we get $[\tau_{1}-\tau_{2}]_{C^{\prime}}=0$, which means that $\eta^{\prime}$ is injective.
Thus, by what we have just proven and by 4.7 the section $\sigma_{C}$ determines a cohomology class $[\sigma_{C}]_{C^{\prime}}$ in $H^{1}(\pi_{1}(C^{\prime},s^{\prime}),\mathbb{Z}^{2g})\cong H^{1}(C^{\prime},\Lambda_{C^{\prime}})$. Again thanks to 4.7, we have that there exists a (possibly transcendental) section $\sigma^{\prime}:C^{\prime}\rightarrow\mathcal{A}^{\prime}|_{C^{\prime}}$ whose cohomology class $[\sigma^{\prime}]$ in $H^{1}(\pi_{1}(C^{\prime},s^{\prime}),\mathbb{Z}^{2g})$ coincides with the class $[\sigma_{C}]_{C^{\prime}}$ defined above, i.e. $[\sigma^{\prime}]=[\sigma_{C}]_{C^{\prime}}$. Therefore, by 3.11 we obtain a section of the form $\sigma_{C}-q_{C}^{*}(\sigma^{\prime})\in\Sigma^{\operatorname{an}}(C)$ that admits a logarithm; let us write
@@
Now, by using properties of the operators $q_{C}^{*}$ and $\operatorname{Tr}$, we will get a contradiction. Applying first the trace operator and then the pullback operator, by part $(ii)$ of 4.5 we obtain:
@@
By part $(vi)$ of 4.5, the section $q^{*}\operatorname{Tr}(\sigma)-\deg{q}\cdot\sigma$ is also an algebraic section admitting a global logarithm on $S(\mathbb{C})$. By 3.10 we must have $\deg{q}\cdot\sigma=q^{*}\operatorname{Tr}(\sigma)$. By 2.1, the section $\deg{q}\cdot\sigma$ is unramified. By 2.3 we conclude that $\sigma$ is an unramified section.
@@ 4.3 Rank of relative monodromy
4.4 The case of thin Monodromy of periods
In [3], André employs Hodge-theoretic methods to study the problems considered in this paper. Under the following hypotheses, he establishes $M_{\sigma}^{\textrm{rel}}\cong\mathbb{Z}^{2g}$:
  • $(i)$
    $\mathcal{A}\to S$ is without fixed part, even after pull-back to a finite étale cover of $S$.
  • $(ii)$
    $\mathbb{Z}\sigma(S)$ is Zariski dense in $\mathcal{A}$.
  • $(iii)$
    The monodromy group $\textrm{Mon}(\mathcal{A})$ is not thin.
Essentially, André’s hypothesis $(iii)$ serves as an alternative to our assumption on the modular map $p$. The geometric distinction is clear: while we require a certain “sparseness” condition on isomorphic fibers of the family, hypothesis $(iii)$ intuitively imposes that the image of the modular map is “sufficiently large.” A more precise comparison of these conditions, or a characterization of when they coincide, would be an interesting direction for further investigation.
Notice that our hypotheses in 1.3 don’t exclude the case when $\textrm{Mon}(\mathcal{A})$ is thin. Indeed, as a consequence of 1.3 and the study of fiber products, we provide below an explicit construction showing a case where $M_{\sigma}^{\textrm{rel}}$ is of maximal rank even though $\textrm{Mon}(\mathcal{A})$ is thin.
Example 4.9.
We consider the same construction as in 2.4. Note that the two elliptic schemes $\mathcal{E}_{1}\to B$ and $\mathcal{E}_{2}\to B$ are not isogenous, since they have a different set of bad reduction. The fundamental group $G$ of $B(\mathbb{C})$ is free on three generators, and the monodromy representation on the periods of $\phi:\mathcal{A}\to B$ gives a homomorphism $\rho:G\to\Gamma_{2}\times\Gamma_{2}$, where
$\Gamma_{2}:=\left\langle\begin{pmatrix}1&2\\ 0&1\end{pmatrix},\;\begin{pmatrix}1&0\\ 2&1\end{pmatrix}\right\rangle\leq\textrm{SL}_{2}(\mathbb{Z}).$
We denote the monodromy group of periods as $\textrm{Mon}(\mathcal{A}):=\rho(G)$. Since $\phi:\mathcal{A}\to B$ is the product of two non-isogenous elliptic schemes, then the Zariski closure of $\textrm{Mon}(\mathcal{A})$ is the whole $\textrm{SL}_{2}\times\textrm{SL}_{2}$ [12, Theorem 5.3.10].
The abelian scheme $\phi\colon\mathcal{A}\to B$ has trivial Mordell-Weil rank. Let us consider the map $\sigma:B\to\mathcal{A}$ defined by
This map is a multisection of the abelian scheme $\phi$, but it becomes a section whether we consider the base change induced by the degree $4$ field extension $\mathbb{C}(B)\subseteq\mathbb{C}(B)(\psi,\mu)$ where $\psi^{2}+\mu^{2}=3$ and $\mu^{2}=1+\lambda$. Denote by $\phi^{\prime}:\mathcal{A}^{\prime}\to B^{\prime}$ the base change we just described. The monodromy of periods $\textrm{Mon}(\mathcal{A}^{\prime})$ is a finite index subgroup of $\textrm{Mon}(\mathcal{A})$: in particular, it is Zariski dense in $\textrm{SL}_{2}\times\textrm{SL}_{2}$ and it has infinite index in $\Gamma_{2}\times\Gamma_{2}$; so $\textrm{Mon}(\mathcal{A}^{\prime})$ is still thin. The morphism $\sigma$ is a section of $\phi^{\prime}$ whose image is not contained in any proper group subscheme, thus the relative monodromy group $M_{\sigma}^{\textrm{rel}}$ is isomorphic to $\mathbb{Z}^{4}$ by [41, Theorem 4.11]. Notice that $\sigma$ is ramified, since the field extension defining $B^{\prime}$ induces a ramified morphism onto $B$ and therefore this gives rise to a ramified (finite flat) modular map.
5 Some applications
We present some useful applications of our results. As a consequence of the non-triviality of relative monodromy we obtain a strong version of Manin’s kernel theorem under some hypothesis on unramified sections. In the case of elliptic schemes over a curve the same result was obtained in [13]. We refer to the Betti coordinates defined in Section 2.2 (see Equation 7).
Theorem 5.1.
Let $\phi:\mathcal{A}\to S$ be an abelian scheme without fixed part such that the modular map $p:S\to T\subseteq\mathbb{A}_{g}$ is generically finite. Denote by $U_{p}\subseteq S$ the Zariski open subset on which $p$ is finite and flat. Assume that all the sections unramified over $U_{p}$ are torsion. Let $\sigma:S\to\mathcal{A}$ be a section. If the Betti coordinates $\beta_{\sigma,i}$ with $i=1,\ldots,2g$ are globally defined functions on $S(\mathbb{C})$, then $\sigma$ is a torsion section. In that case, the Betti coordinates are rational constants.
@@ 5 Some applications
If the Betti coordinates are globally defined on $S(\mathbb{C})$ then the relative monodromy group $M_{\sigma}^{\textrm{rel}}$ is trivial. The claim then follows by 1.2.
Our theorem about the rank of the relative monodromy group 1.3 gives a different proof of a result due to André [2, Theorem 3] about the algebraic independence of logarithms under our hypothesis on the abelian scheme $\mathcal{A}\to S$. Let’s briefly describe the setting.
Recall that the monodromy action of the fundamental group $\pi_{1}(S(\mathbb{C}),s)$ on periods and the logarithm of a section $\sigma$ induces a representation $\theta_{\sigma}:\pi_{1}(S(\mathbb{C}),s)\rightarrow\textrm{SL}_{2g+1}(\mathbb{Z})$ described in Equation 16. Consider the projection $\theta_{\sigma}(\pi_{1}(S(\mathbb{C}),s))\to\textrm{Sp}_{2g}(\mathbb{Z})$, whose image is the monodromy group $\textrm{Mon}(\mathcal{A})$. By general theory, the Zariski closure $\overline{\theta_{\sigma}(\pi_{1}(S(\mathbb{C}),s))}$ in $\textrm{SL}_{2g+1}(\mathbb{Z})$ is the differential Galois group of the Picard-Vessiot extension of $\mathbb{C}(S)$ generated by the coordinates of $\omega_{1},\ldots,\omega_{2g},\log_{\sigma}$ and their directional derivatives along a tangent vector field $\partial$ with respect to the Gauss-Manin connection. Clearly the homomorphism $\theta_{\sigma}(\pi_{1}(S(\mathbb{C}),s))\to\textrm{Sp}_{2g}(\mathbb{Z})$ extends to an algebraic group homomorphism
$\xi:\overline{\theta_{\sigma}(\pi_{1}(S(\mathbb{C}),s))}\to\textrm{Sp}_{2g}(\mathbb{Z})$
@@
Let $\phi:\mathcal{A}\to S$ be an abelian scheme without fixed part such that the modular map $p:S\to T\subseteq\mathbb{A}_{g}$ is generically finite. Denote by $U_{p}\subseteq S$ the Zariski open subset on which $p$ is finite and flat. Assume that the action of the monodromy group $\textrm{Mon}(\mathcal{A})$ is irreducible. If a section $\sigma:S\to\mathcal{A}$ is ramified over $U_{p}$, then the kernel $\ker{\xi}$ is isomorphic to $\mathbb{G}_{a}^{2g}$. In particular, the coordinates of $\log_{\sigma}$ have transcendence degree $g$ over $\mathbb{C}(S)(\omega_{1},\ldots,\omega_{2g})$, where $\mathbb{C}(S)(\omega_{1},\ldots,\omega_{2g})$ denotes the extension of $\mathbb{C}(S)$ generated by the coordinates of periods.